# Generating a Paracosm for Training-Free Zero-Shot Composed Image Retrieval

### v1

@@ 1 Introduction
Composed Image Retrieval (CIR) formulates new scenarios in internet search [10] and e-commerce [39, 54], where it retrieves a target image from a database based on a user-provided multimodal query, which consists of a reference image and a modification text [51]. The text describes how to alter the reference image for which CIR algorithms should find the best matched image from the database [29, 1, 50].
@@ 2 Related Work
Zero-Shot Composed Image Retrieval (ZS-CIR) aims to solve CIR without directly training on triplet data [36]. It does not mean developing training-free methods but allows training in an indirect manner [36, 4, 43, 16, 59, 42]. For example, many methods train a textual inversion network on existing image datasets [38, 35], mapping the reference image to a pseudo-word token, which, along with the modification text, is used in cross-modal matching with dataset images. A couple of recent methods [52, 26] use pretrained generative models [34, 37, 62] to synthesize a synthetic image (similar to our “mental image”), termed as pseudo-targe images, for a given multimodal query and use it to augment the textual feature (i.e., output by the textual inverse network) to compute similarity scores with dataset images. Importantly, training-free ZS-CIR methods [20, 57, 44] have emerged that leverage an LMM to generate a description for a given multimodal query. This avoids training a separate textual inversion network. Notably, existing ZS-CIR methods have not considered generating either descriptions or synthetic images for database images. In contrast, our work does so, motivated to address ZS-CIR from first principles. Specifically, we propose to generate the “mental image” for the query and synthetic counterparts for real dataset images, mitigating synthetic-to-real domain gaps for better performance.
Foundation Models (FMs), pretrained on various formats of web-scale data, demonstrate unprecedented performance on downstream tasks in a zero-shot manner. FMs are extensively exploited in the CIR literature [25, 6, 56, 28, 7]. First, Vision-Language Models (VLMs), pretrained on web-scale image-text pairs [33, 18], are commonly used in CIR to extract features of modification texts and images, enabling cross-modality matching [20, 44, 5, 26]. Second, Large Language Models (LLMs), pretrained on massive text corpus [41, 8, 48, 2], are used in CIR for generating a target image description by incorporating the reference image caption and the modification text [57, 26]. Third, Large Multimodal Models (LMMs) [32, 53, 45, 3], pretrained on web-scale multimodal data and typically larger than VLMs in parameter size, are used in CIR to generate image captions [20, 57, 26] for multimodal queries. Earlier CIR works use LMMs to create triplet data for supervised training [28, 59, 42]. Most training-free ZS-CIR methods [20, 44, 57] use LMMs to generate textual descriptions for multimodal queries and utilize a VLM to match the descriptions with database images, transforming the CIR task to a text-to-image retrieval problem. Following this literature, we extensively exploit LMMs to generate “mental images” for multimodal queries and synthetic counterparts for real database images. Matching them effectively mitigates synthetic-to-real domain gaps and boosts CIR performance.
@@ 3.2 The Proposed Method: Paracosm
Processing Multimodal Query. For a multimodal query, Paracosm first generates the “mental image” ${\bf I}_{mental}$ using an LMM [53]. Based on this mental image ${\bf I}_{mental}$, it further generates a brief description ${\mathbf{t}}_{query}$, which can be thought of as the description of the target image ${\bf I}_{target}$. Some recent methods [20, 44, 57] propose to generate such a description based on the multimodal query, e.g., by first generating a description for the reference image and then revising it with the modification using an LLM. Fig.˜3 shows that solely relying on the generated description misses crucial visual information, whereas the mental image keeps rich information and incorporating it boosts CIR. Table˜5 quantitatively verifies this.
Processing Database Images. As the generated mental images are synthetic and have synthetic-to-real domain gaps, directly matching them with real images from the database is suboptimal. To mitigate this gap, Paracosm also generates a synthetic visual ${\bf I}_{syn}^{i}$ for each database image ${\bf I}^{i}$. To this end, Paracosm first leverages an LMM [46], to generate a detailed description. It then uses this description as prompt to a text-to-image generation model [53], generating a synthetic counterpart for each database image. The real database images and synthetic counterparts are used jointly for matching a multimodal query. Table˜5 validates this design choice.
@@ 3.3 Remarks
Paracosm makes better use of LMMs. Existing ZS-CIR approaches [57, 44, 20] also leverage LMMs but focus on generating descriptions for multimodal queries and transforming the CIR problem into a text-to-image retrieval problem. However, text descriptions alone cannot sufficiently capture rich visual information crucial to CIR and hence often lead to incorrect retrieved images, as demonstrated in Fig.˜3. A couple of recent works [26, 52] realize the importance of incorporating pseudo-targe images to facilitate CIR, e.g., by exploiting a pretrained text-to-image (T2I) generative model [62] to synthesize a pseudo-target image based on a description. Differently, we leverage an LMM that has image editing capability, i.e., editing the reference image based on the modification text into a “mental image”. As shown in Table˜4, editing reference images yields better CIR results than T2I generation of pseudo-target images. Moreover, Paracosm exploits an LMM to generate detailed descriptions for each database image, and uses this description as prompt to generate a synthetic counterpart.
@@
Computation cost. As Paracosm extensively exploits LMMs, it has a high computation cost, especially for generating images, i.e., the mental images for the query and synthetic counterparts of database images. Yet, this is not a flaw pertaining only to Paracosm, as existing works also rely on generative models for retrieval. For example, prevailing ZS-CIR methods use LMMs to generate descriptions [57, 44, 20, 26], and some turn to generative models to synthesize images [26, 52]. Importantly, like these methods, Paracosm can process database images ahead of time, and does not require computing features for them on the fly during inference. So the major computation overhead is on the process of a given multimodal query in inference, including generating the corresponding mental images and their descriptions. It is worth noting that efficient inference in generative models is an important topic and has been greatly advanced through model optimization [58, 19] and optimized implementation [13]. Therefore, Paracosm should not suffer from computation in the long run.
@@
OpenCLIP OSrCIR w/ GPT-4o reproduced 32.96 63.28 74.99 92.29 66.72 84.39 93.08 18.77 19.33 21.04 22.02
ViT-B/32 OSrCIR w/ Qwen2.5-VL reproduced 29.18 58.43 70.27 89.69 62.17 80.58 90.46 14.47 15.23 16.70 17.55
@@
OpenCLIP OSrCIR w/ GPT-4o reproduced 35.47 65.06 76.96 92.99 67.42 85.66 93.30 22.75 23.51 25.73 26.78
ViT-L/14 OSrCIR w/ Qwen2.5-VL reproduced 31.59 61.06 72.60 90.84 63.35 81.86 90.82 17.88 18.72 20.51 21.39
@@ 4 Experiments
OpenCLIP OSrCIR w/ GPT-4o reproduced 31.26 51.08 28.36 48.98 34.63 56.71 31.41 52.26
ViT-B/32 OSrCIR w/ Qwen2.5-VL reproduced 27.97 47.25 23.15 44.62 28.51 50.18 26.54 47.35
@@
OpenCLIP OSrCIR w/ GPT-4o reproduced 31.40 50.74 28.61 49.43 34.63 55.53 31.55 51.90
ViT-L/14 OSrCIR w/ Qwen2.5-VL reproduced 28.75 46.32 24.74 43.98 30.04 51.20 27.84 47.16
@@
  • Training-free methods. CIReVL [20] uses an LMM to generate a description of a given reference image, and then an LLM to modify this description based on the modification text. It uses the modified description for cross-modality retrieval in the image database. LDRE [57] proposes to generate multiple diverse descriptions to improve CIR performance. OSrCIR [44] carefully designs the prompt to an LMM with Reflective Chain-of-Thought to improve the output description quality for a multimodal query.
@@ 4.1 Benchmarking Results
OpenCLIP OSrCIR w/ GPT-4o reproduced 14.2 25.9 36.5
ViT-B/32 OSrCIR w/ Qwen2.5-VL reproduced 15.5 26.7 36.6
@@
OpenCLIP OSrCIR w/ GPT-4o reproduced 15.0 25.9 35.7
ViT-L/14 OSrCIR w/ Qwen2.5-VL reproduced 15.5 27.0 36.9
@@ 4.2 Analyses and Ablation Study
$\lambda$ for combining visual and textual features. Paracosm (Fig.˜2) employs a hyperparameter $\lambda$ (Eq. 1) to control the contribution of modification texts from the query. Fig.˜4 studies its effects on the final CIR performance over the four benchmarks. Interestingly, $\lambda=0.3$ consistently produces the highest numeric metrics on all the benchmarks. We set $\lambda=0.3$ throughout our work.
Reference image editing vs. Text-to-Image generation. When generating mental images for multimodal queries, our Paracosm leverages the LMM to directly edit the reference images based on modification texts. In comparison, another workaround is T2I generation, as done in the recent work [26]: generating descriptions for reference images using an LMM, then modifying the descriptions based on modification texts using an LLM, and lastly generating mental images based on the modified descriptions. Table˜4 compares them, clearly showing that Paracosm’s design choice leads to better performance.
Figure 4: Analysis of $\lambda$ which controls the importance of incorporating modification text in Eq. 1. Interestingly, on all datasets, setting $\lambda=0.3$ consistently yields the highest numeric metrics reported for all the datasets.
Ablation Study. We ablate core components in Paracosm with the CLIP ViT-B/32 VLM on CIRR and CIRCO. Table˜5 excerpts salient results; supplementary Section˜F provides comprehensive details.
@@
Table 4: Comparing different generation methods for the mental images within Paracosm. Here, we adopt the CLIP ViT-B/32 VLM for Paracosm and report results on the CIRR test set. “T2I Generation” means first generating pseudo-target descriptions and then basing on them to generate mental images. “Image Edit” means our design choice for generating mental images in Paracosm, i.e., directly editing reference images based on modification texts. Clearly, the latter performs better.
@@
Image Edit 32.27 62.60 75.16 92.60
@@ A Additional Details of Benchmark Datasets
In this section, we provide additional details of four benchmark datasets, i.e., CIRR [29], CIRCO [1], Fashion IQ [54], and GeneCIS [50]. In addition, we summarize the datasets used in our experiments in Table˜6.
@@ D Further Analyses of LMMs
We further analyze the impact of using different LMMs for generating textual descriptions of mental images. Specifically, we replace Qwen2.5-VL-7B-Instruct [46] used in Paracosm with GPT-4o, the LMM employed by OSrCIR [44]. As shown in Tab.˜8, switching to GPT-4o yields only marginal improvements in retrieval performance. Notably, the best performance across all datasets is achieved when $\lambda$ = 0.3 (Fig.˜4), indicating that the final retrieval performance is relatively insensitive to the choice of LMM for mental image description generation. This indicates that the core strength of Paracosm lies not in the scale of LMMs but in the “paracosm” constructed by our method. Additionally, we generate brief textual descriptions ($\mathbf{t}^{i}_{brief}$) for candidate images and incorporate them into Paracosm, as illustrated in Fig.˜9. However, as shown in Tab.˜8, this strategy yields performance gains only on the CIRR and CIRCO datasets. Consequently, we opt to exclude $\mathbf{t}^{i}_{brief}$ from our final pipeline.

### v2

@@ 1 Introduction
Composed Image Retrieval (CIR) formulates new scenarios in internet search [10, 24] and e-commerce [43, 61], where it retrieves a target image from a database based on a user-provided multimodal query, which consists of a reference image and a modification text [58]. The text describes how to alter the reference image for which CIR algorithms should find the best matched image from the database [33, 1, 57].
@@ 2 Related Work
Zero-Shot Composed Image Retrieval (ZS-CIR) aims to solve CIR without directly training on triplet data [40]. It does not mean developing training-free methods but allows training in an indirect manner [40, 4, 49, 18, 66, 48]. For example, many methods train a textual inversion network on existing image datasets [42, 39], mapping the reference image to a pseudo-word token, which, along with the modification text, is used in cross-modal matching with dataset images. A couple of recent methods [59, 30] use pretrained generative models [38, 41, 70] to synthesize a synthetic image (similar to our “mental image”), termed as pseudo-targe images, for a given multimodal query and use it to augment the textual feature (i.e., output by the textual inverse network) to compute similarity scores with dataset images. Importantly, training-free ZS-CIR methods [23, 64, 50, 45, 12] have emerged that leverage an LMM to generate a description for a given multimodal query. This avoids training a separate textual inversion network. Notably, existing ZS-CIR methods have not considered generating either descriptions or synthetic images for database images. In contrast, our work does so, motivated to address ZS-CIR from first principles. Specifically, we propose to generate the “mental image” for the query and synthetic counterparts for real dataset images, mitigating synthetic-to-real domain gaps for better performance.
Foundation Models (FMs), pretrained on various formats of web-scale data, demonstrate unprecedented performance on downstream tasks in a zero-shot manner. FMs are extensively exploited in the CIR literature [29, 6, 63, 32, 7]. First, Vision-Language Models (VLMs), pretrained on web-scale image-text pairs [37, 21], are commonly used in CIR to extract features of modification texts and images, enabling cross-modality matching [46, 23, 50, 5, 30]. Second, Large Language Models (LLMs), pretrained on massive text corpus [47, 8, 55, 2], are used in CIR for generating a target image description by incorporating the reference image caption and the modification text [64, 30]. Third, Large Multimodal Models (LMMs) [36, 60, 51, 3], pretrained on web-scale multimodal data and typically larger than VLMs in parameter size, are used in CIR to generate image captions [23, 64, 30] for multimodal queries. Earlier CIR works use LMMs to create triplet data for supervised training [32, 66, 48]. Most training-free ZS-CIR methods [23, 50, 64] use LMMs to generate textual descriptions for multimodal queries and utilize a VLM to match the descriptions with database images, transforming the CIR task to a text-to-image retrieval problem. Following this literature, we extensively exploit LMMs to generate “mental images” for multimodal queries and synthetic counterparts for real database images. Matching them effectively mitigates synthetic-to-real domain gaps and boosts CIR performance.
@@ 3.2 The Proposed Method: Paracosm
Processing Multimodal Query. For a multimodal query, Paracosm first generates the “mental image” ${\bf I}_{mental}$ using an LMM [60]. Based on this mental image ${\bf I}_{mental}$, it further generates a brief description ${\mathbf{t}}_{query}$, which can be thought of as the description of the target image ${\bf I}_{target}$. Some recent methods [23, 50, 64] propose to generate such a description based on the multimodal query, e.g., by first generating a description for the reference image and then revising it with the modification using an LLM. Fig.˜3 shows that solely relying on the generated description misses crucial visual information, whereas the mental image keeps rich information and incorporating it boosts CIR. Table˜6 quantitatively verifies this.
Processing Database Images. As the generated mental images are synthetic and have synthetic-to-real domain gaps, directly matching them with real images from the database is suboptimal. One strategy to mitigate this gap is to create synthetic counterparts for database images and use such to assist matching [69, 15, 19]. Specifically, Paracosm first leverages an LMM [53], to generate a detailed description for each database image. It then uses this description as prompt to a text-to-image generation model [60], generating a synthetic counterpart ${\bf I}_{syn}^{i}$ for each database image ${\bf I}^{i}$. The real database images and synthetic counterparts are used jointly for matching a multimodal query. Table˜6 validates this design choice.
@@ 3.3 Remarks
Paracosm makes better use of LMMs. Existing ZS-CIR approaches [64, 50, 23] also leverage LMMs but focus on generating descriptions for multimodal queries and transforming the CIR problem into a text-to-image retrieval problem. However, text descriptions alone cannot sufficiently capture rich visual information crucial to CIR and hence often lead to incorrect retrieved images, as demonstrated in Fig.˜3. A couple of recent works [30, 59] realize the importance of incorporating pseudo-targe images to facilitate CIR, e.g., by exploiting a pretrained text-to-image (T2I) generative model [70] to synthesize a pseudo-target image based on a description. Differently, we leverage an LMM that has image editing capability, i.e., editing the reference image based on the modification text into a “mental image”. As shown in Table˜5, editing reference images yields better CIR results than T2I generation of pseudo-target images. Moreover, Paracosm exploits an LMM to generate detailed descriptions for each database image, and uses this description as prompt to generate a synthetic counterpart.
@@
Computation cost. As Paracosm extensively exploits LMMs, it has a high computation cost, especially for generating images, i.e., the mental images for the query and synthetic counterparts of database images. Yet, this is not a flaw pertaining only to Paracosm, as existing works also rely on generative models for retrieval. For example, prevailing ZS-CIR methods use LMMs to generate descriptions [64, 50, 23, 30], and some turn to generative models to synthesize images [30, 59]. Importantly, like these methods, Paracosm can process database images ahead of time, and does not require computing features for them on the fly during inference. So the major computation overhead is on the process of a given multimodal query in inference, including generating the corresponding mental images and their descriptions. As efficient inference in generative models is an important topic and has been greatly advanced through model optimization [65, 22] and optimized implementation [14], Paracosm would not suffer from computation in the long run.
@@
OpenCLIP AutoCIR [12] w/ GPT-4o-mini KDD’25 30.53 59.42 72.19 91.47 65.11 84.02 92.70 18.82 19.41 21.38 22.32
ViT-B/32 CoTMR [45] w/ Qwen2-VL ICCV’25 31.50 60.80 73.04 91.06 66.61 84.50 92.55 22.23 22.78 24.68 25.74
OSrCIR w/ GPT-4o reproduced 32.96 63.28 74.99 92.29 66.72 84.39 93.08 18.77 19.33 21.04 22.02
OSrCIR w/ Qwen2.5-VL reproduced 29.18 58.43 70.27 89.69 62.17 80.58 90.46 14.47 15.23 16.70 17.55
@@
OpenCLIP AutoCIR [12] w/ GPT-4o-mini KDD’25 31.81 61.95 73.86 92.07 67.21 84.89 93.13 24.05 25.14 27.35 28.36
ViT-L/14 CoTMR [45] w/ Qwen2-VL ICCV’25 35.02 64.75 76.18 92.51 69.39 85.75 93.33 27.61 28.22 30.61 31.70
OSrCIR w/ GPT-4o reproduced 35.47 65.06 76.96 92.99 67.42 85.66 93.30 22.75 23.51 25.73 26.78
OSrCIR w/ Qwen2.5-VL reproduced 31.59 61.06 72.60 90.84 63.35 81.86 90.82 17.88 18.72 20.51 21.39
@@ 4 Experiments
OpenCLIP AutoCIR [12] w/ GPT-4o-mini KDD’25 32.43 51.67 26.52 46.36 33.96 56.09 30.97 51.37
ViT-B/32 CoTMR [45] w/ Qwen2-VL ICCV’25 33.42 53.93 31.09 54.54 38.40 61.14 34.30 56.54
OSrCIR w/ GPT-4o reproduced 31.26 51.08 28.36 48.98 34.63 56.71 31.41 52.26
OSrCIR w/ Qwen2.5-VL reproduced 27.97 47.25 23.15 44.62 28.51 50.18 26.54 47.35
@@
OpenCLIP AutoCIR [12] w/ GPT-4o-mini KDD’25 34.00 53.43 24.94 45.81 33.10 55.58 30.68 51.60
ViT-L/14 CoTMR [45] w/ Qwen2-VL ICCV’25 35.43 54.91 31.18 55.04 38.55 61.33 35.05 57.09
OSrCIR w/ GPT-4o reproduced 31.40 50.74 28.61 49.43 34.63 55.53 31.55 51.90
OSrCIR w/ Qwen2.5-VL reproduced 28.75 46.32 24.74 43.98 30.04 51.20 27.84 47.16
@@
  • Training-free methods. CIReVL [23] uses an LMM to generate a description of a given reference image, and then an LLM to modify this description based on the modification text. It uses the modified description for cross-modality retrieval in the image database. LDRE [64] proposes to generate multiple diverse descriptions to improve CIR performance. AutoCIR [12] employs automatic multi-agent collaboration to iteratively refine retrieval queries through self-correction. CoTMR [45] leverages an LMM to generate both global captions and fine-grained object-level descriptions combined with a multi-grained scoring that significantly improves the retrieval performance. OSrCIR [50] carefully designs the prompt to an LMM with Reflective Chain-of-Thought to improve the output description quality for a multimodal query.
@@ 4.1 Benchmarking Results
OpenCLIP CoTMR [45] w/ Qwen2-VL ICCV’25 17.9 29.6 39.5
ViT-B/32 OSrCIR w/ GPT-4o reproduced 14.2 25.9 36.5
OSrCIR w/ Qwen2.5-VL reproduced 15.5 26.7 36.6
@@
OpenCLIP CoTMR [45] w/ Qwen2-VL ICCV’25 17.6 29.6 39.8
ViT-L/14 OSrCIR w/ GPT-4o reproduced 15.0 25.9 35.7
OSrCIR w/ Qwen2.5-VL reproduced 15.5 27.0 36.9
@@
Efficiency Comparisons. For a comprehensive analysis, we evaluate costs of both Paracosm and some recent methods under the same experimental condition in  Table˜4. AutoCIR [12] and OSrCIR [50] use GPT [36], whereas CoTMR [45] and our Paracosm use Qwen [3]. We run all the methods on an NVIDIA A100 GPU with OpenCLIP ViT-L/14 for feature extraction. Computational costs are measured on the CIRCO benchmark, comprising >>123K database images. The results show that (1) Paracosm incurs substantial GPU hours to process database images, (2) but its inference efficiency is comparable to AutoCIR when both use API, and (3) importantly, Paracosm significantly outperforms these methods!
Table 4: Efficiency comparisons. Computational costs are measured on the CIRCO benchmark. We break down the costs into (1) database processing (w.r.t GPU hours, and storage for synthetic images and features), and (2) per-query inference (w.r.t memory consumption and latency). Notably, synthetic database images are precomputed and not required to store.
Method GPU storage in GB inference
hrs img. feat. (GB) (sec)
AutoCIR (KDD’25) [12] 0.3 n/a 0.38 2.3 24
CoTMR (ICCV’25) [45] 0.3 n/a 0.38 70.2 1
OSrCIR (CVPR’25) [50] 0.3 n/a 0.38 2.3 3
Paracosm 205.7 41.4 0.38 2.7 14
@@ 4.2 Analyses and Ablation Study
$\lambda$ for combining visual and textual features. Paracosm (Fig.˜2) employs a hyperparameter $\lambda$ (Eq. 1) to control the contribution of modification texts from the query. We tuned $\lambda$ on the CIRR validation set using CLIP ViT-B/32 and set $\lambda=0.3$ throughout our work. Fig.˜4 studies its effects on the final CIR performance over the four benchmarks. Interestingly, $\lambda=0.3$ consistently produces the highest numeric metrics on all the benchmarks.
Reference image editing vs. Text-to-Image generation. When generating mental images for multimodal queries, our Paracosm leverages the LMM to directly edit the reference images based on modification texts. In comparison, another workaround is T2I generation, as done in the recent work [30]: generating descriptions for reference images using an LMM, then modifying the descriptions based on modification texts using an LLM, and lastly generating mental images based on the modified descriptions. Table˜5 compares them, clearly showing that Paracosm’s design choice leads to better performance.
Different image generator. We compare the results of Paracosm using Qwen-Image-Edit [60] and another open-source image generator LongCat-Image-Edit [52] on the CIRR test set without tuning hyperparameters or prompts. The results in Table˜5 show stable performance of Paracosm with different generators.
Figure 4: Analysis of $\lambda$ which controls the importance of incorporating modification text in Eq. 1. We set $\lambda=0.3$ based on the results of the CIRR validation set. Interestingly, on all datasets, setting $\lambda=0.3$ consistently yields the highest numeric metrics reported for all the datasets.
Ablation Study. We ablate core components in Paracosm with the CLIP ViT-B/32 VLM on CIRR and CIRCO. Table˜6 excerpts salient results; supplementary Section˜F provides comprehensive details.
@@
Table 5: Comparing different generation methods for the mental images within Paracosm. Here, we adopt the CLIP ViT-B/32 VLM for Paracosm and report results on the CIRR test set. “T2I Generation” means first generating pseudo-target descriptions and then basing on them to generate mental images. “Image Edit” means our design choice for generating mental images in Paracosm, i.e., directly editing reference images based on modification texts. Clearly, the latter performs better. Moreover, for “Image Edit”, we compare using Qwen-Image-Edit vs. LonaCat-Image-Edit as image generator. The results show stable performance of Paracosm with different generators.
@@
Image Edit w/ Qwen 32.27 62.60 75.16 92.60
Image Edit w/ LongCat 32.12 62.20 74.43 92.24
@@ A Additional Details of Benchmark Datasets
In this section, we provide additional details of four benchmark datasets, i.e., CIRR [33], CIRCO [1], Fashion IQ [61], and GeneCIS [57]. In addition, we summarize the datasets used in our experiments in Table˜7.
@@ D Further Analyses of LMMs
We further analyze the impact of using different LMMs for generating textual descriptions of mental images. Specifically, we replace Qwen2.5-VL-7B-Instruct [53] used in Paracosm with GPT-4o, the LMM employed by OSrCIR [50]. As shown in Tab.˜9, switching to GPT-4o yields only marginal improvements in retrieval performance. Notably, the best performance across all datasets is achieved when $\lambda$ = 0.3 (Fig.˜4), indicating that the final retrieval performance is relatively insensitive to the choice of LMM for mental image description generation. This indicates that the core strength of Paracosm lies not in the scale of LMMs but in the “paracosm” constructed by our method. Additionally, we generate brief textual descriptions ($\mathbf{t}^{i}_{brief}$) for candidate images and incorporate them into Paracosm, as illustrated in Fig.˜9. However, as shown in Tab.˜9, this strategy yields performance gains only on the CIRR and CIRCO datasets. Consequently, we opt to exclude $\mathbf{t}^{i}_{brief}$ from our final pipeline.